OpenStudy (anonymous):

please help. i get two diff ans. If the interest rate is 3% and a total of $4,370.91 will be paid to you at the end of 3 years, what is the present value of that sum? A) $3900, B) $3947 C) $3,977.40 D) $4000

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (wolf1728):

You want to know the amount you started with so the formula is: Principal = Total ÷ (1 + Rate)^years Principal = 4,370.91 ÷ (1.03)^3 Principal = 4,370.91 ÷ 1.092727 Principal = $4,000.00 Calculator to test your answer http://www.1728.org/compint.htm

OpenStudy (anonymous):

that is exactly what i did and got $4000, but when i tried to check my answer, by google, i got $3900. i got $4000 four diff times. THANK U SO MUCH!

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

u get a medal

OpenStudy (anonymous):

thanks for the website.